# Enhancing Variable Selection in Large-scale Logistic Regression: Leveraging Manual Labeling with Beneficial Noise

### v1

@@ IV-B Parallel ADMM for ALASSO-PLR with Manual Labels
$\bm{X}=(\bm{X}_{1}^{\top},\bm{X}_{2}^{\top},\dots,\bm{X}_{G}^{\top})^{\top}$ (32)
and $\bm{S}=(\bm{S}_{1}^{\top},\bm{S}_{2}^{\top},\dots,\bm{S}_{G}^{\top})^{\top}$
are distributedly stored across $G$ local sub-machines. Here $\bm{S}_{g}\in\mathbb{R}^{n_{g}}$, $\bm{X}_{g}\in\mathbb{R}^{n_{g}\times p}$ and $\sum\limits_{g=1}^{G}n_{g}=n$. The central machine takes in information from local sub-machines, consolidates it, and sends it back to relevant locals. The decomposition in (LABEL:mdata) makes the algorithm naturally parallel across multiple local sub-machines. Each local sub-machine can independently solve its allotted subproblem. Then, the central and local sub-machines communicate and coordinate solutions via variable updates. This distributed storage allows for parallel processing and efficient handling of large-scale data. For a more detailed discussion, please refer to Chapters 7 and 8 in [31].
@@
$\min_{\bm{\beta},\bm{r}_{g}}$ $\sum_{g=1}^{G}\sum_{i=1}^{n_{g}}\Big{[}-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\Big{]}$
$+\lambda\|\bm{w}\odot\bm{\beta}\|_{1},$
s.t. $\bm{r}_{g}=\bm{X}_{g}\bm{\beta},\ g=1,2,\dots,G.$ (33)
It is not difficult to verify that (IV) and (IV-B) are equivalent. The augmented Lagrangian of (IV-B) is given by
$L_{\mu}(\bm{\beta},\bm{r}_{g},\bm{u}_{g})$ $=\sum_{g=1}^{G}\sum_{i=1}^{n_{g}}\Big{[}-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\Big{]}$
$+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\sum_{g=1}^{G}\bm{u}_{g}^{\top}(\bm{X}_{g}\bm{\beta}-\bm{r}_{g})$
$+\frac{\mu}{2}\sum\limits_{g=1}^{G}\|\bm{X}_{g}\bm{\beta}_{g}-\bm{r}_{g}\|_{2}^{2},$ (34)
where $\bm{u}_{g}\in\mathbb{R}^{n_{g}}$ is the dual variable and $(\bm{u}_{1}^{\top},\bm{u}_{2}^{\top},\dots,\bm{u}_{G}^{\top})^{\top}=\bm{u}$. Similar to (22), the update of the variables in the iterative process (with linearization) can be written as
$\bm{\beta}^{k+1}\leftarrow&\mathop{\arg\min}\limits_{\bm{\beta}}\left\{L_{\mu}(\bm{\beta},\bm{r}_{g}^{k},\bm{u}_{g}^{k})+\frac{1}{2}\|\bm{\beta}-\bm{\beta}^{k}\|^{2}_{{\bm{E}}_{1}}\right\};\\ \bm{r}_{g}^{k+1}\leftarrow&\mathop{\arg\min}\limits_{\bm{r}}\left\{L_{\mu}(\bm{\beta}^{k+1},\bm{r}_{g},\bm{u}_{g}^{k})\right.\\ +&\left.(c_{0}-o(1))\|\bm{r}_{g}-\bm{r}_{g}^{k}\|^{2}_{\bm{I}_{n_{g}}}\right\},g=1,2,\dots,G;\\ \bm{u}_{g}^{k+1}\leftarrow&\bm{u}_{g}^{k}-\mu(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{r}_{g}^{k+1}),g=1,2,\dots,G.\end{aligned}\right.$ (35)
@@
Looking back at the equation in (LABEL:lbeta3), in a distributed environment, the update of $\bm{\beta}$ encounters two issues. One issue is that the complete dataset $\bm{X}$ is unavailable (since $\bm{X}$ is stored in a distributed manner and we can only access $\bm{X}_{g}$). The other issue is the computation of the value of $\bm{\eta}$ based on $\bm{X}$. Based on the suggestion in [39], we can reformulate the update of $\bm{\beta}$ in (25) as
$\bm{\beta}^{k+1}\leftarrow\underset{\bm{\beta}}{\mathrm{argmin}}$$\left\{\lambda\left\lVert\bm{w}\odot\bm{\beta}\right\rVert_{1}\right.$ (36)
$+$ $\left.\frac{\eta}{2}\left\lVert\bm{\beta}-\bm{\beta}^{k}+\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right\rVert_{2}^{2}\right\},$
@@
$\bm{\beta}^{k+1}\leftarrow\text{sign}$ $\left(\bm{\beta}^{k}-\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right)$ (37)
$\odot$ $\left(\left|\bm{\beta}^{k}-\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right|-\frac{\lambda\bm{w}}{\eta}\right)_{+}.$
Therefore, the first issue with the $\bm{\beta}$ update has been resolved. The second issue pertains to the calculation of $\eta$. The calculation of $\eta$ depends on $\bm{X}$. However, each sub-machine only has $\bm{X}_{g}$. According to the suggestion in [39], $\eta_{g}=\sigma(\mu\bm{X}_{g}^{\top}\bm{X}_{g})$ for each sub-machine can be calculated first. Since $\bm{X}^{\top}\bm{X}=\sum\limits_{g=1}^{G}\bm{X}_{g}^{\top}\bm{X}_{g}$, we have $\sum\limits_{g=1}^{G}\eta_{g}\geq\eta$. Therefore, $\sum\limits_{g=1}^{G}\eta_{g}$ can be used to substitute the $\eta$ calculated from the full data. So far, two issues related to the update of $\bm{\beta}$ in distributed environments have been resolved.
Similar to the derivation of the closed form solution shown in (30), we can obtain the update rule for $\bm{r}_{g}$ in a distributed environment,
$\bm{r}_{g}^{k+1}\leftarrow$ $\left[\Phi(\bm{r}_{g}^{k})+\mu\bm{I}_{n_{g}}\right]^{-1}\left[\mu(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{u}_{g}^{k}/\mu)\right.$
$+$ $\left.\Phi(\bm{r}_{g}^{k})\bm{r}_{g}^{k}-\phi^{{}^{\prime}}(\bm{r}_{g}^{k})+\bm{S}_{g}\right],$ (38)
@@
   Central machine:  1. Receive $\eta_{g}$ (only once) and $\bm{\xi}_{g}^{k}$ transmitted by all sub-machines, 2. Update $\bm{\beta}^{k+1}$ using (37),  3. Send $\bm{\beta}^{k+1}$ to the each sub-machines.
   Local machines:    for $g=1,2,\dots,G$ (in parallel)  1. Receive $\bm{\beta}^{k+1}$ transmitted by the central machine,  2. Update $\bm{r}^{k+1}$ using (IV-B),  3. Update $\bm{u}_{g}^{k+1}\leftarrow\bm{u}_{g}^{k}-\mu(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{r}_{g}^{k+1}),$  4. Compute $\bm{\xi}_{g}^{k+1}=\frac{\mu\bm{X}_{g}^{\top}(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{r}_{g}^{k+1}-\bm{u}_{g}^{k+1}/\mu)}{\eta}$,  5. Send $\bm{\xi}^{k+1}_{g}$ to the central machine.
@@
$\left\{\hat{\bm{\beta}}^{k},\hat{\bm{r}}^{k},\hat{\bm{d}}^{k}\right\}=\left\{\bar{\bm{\beta}}^{k},\bar{\bm{r}}^{k},\bar{\bm{d}}^{k}\right\},\quad\text{for all }k.$ (39)
The proof of Theorem 4 can be found in Appendix A-D. The conclusion of Theorem 4 indicates that when using the same initial values, the solutions of Algorithm 1 and Algorithm 2 will be the same, which is what Wu et al. [39] called partition insensitivity. In parallel computing, insensitivity to data partitioning can enhance computational efficiency (by reducing communication overhead and increasing parallelism), improve system stability (by minimizing the risk of load imbalance and enhancing fault tolerance), and simplify algorithm design and implementation (by lowering design complexity and facilitating code maintenance). However, in reality,$\sum\limits_{g=1}^{G}\eta_{g}$ (Algorithm 2) is usually greater than $\eta$ (Algorithm 1) in applications, but this does not affect the convergence of Algorithm 2 with Theorem 3. Simply replace $\eta$ in $\bm{E}_{1}$ with $\sum\limits_{g=1}^{G}\eta_{g}$ in Theorem 3.
V Numerical Experiments

### v2

@@ 4.2 Parallel ADMM for ALASSO-PLR with Manual Labels
$\bm{X}=(\bm{X}_{1}^{\top},\bm{X}_{2}^{\top},\dots,\bm{X}_{G}^{\top})^{\top}\ \ \text{and}\ \ \bm{S}=(\bm{S}_{1}^{\top},\bm{S}_{2}^{\top},\dots,\bm{S}_{G}^{\top})^{\top}$ (4.12)
are distributedly stored across $G$ local sub-machines. Here $\bm{S}_{g}\in\mathbb{R}^{n_{g}}$, $\bm{X}_{g}\in\mathbb{R}^{n_{g}\times p}$ and $\sum\limits_{g=1}^{G}n_{g}=n$. The central machine takes in information from local sub-machines, consolidates it, and sends it back to relevant locals. The decomposition in (4.12) makes the algorithm naturally parallel across multiple local sub-machines. Each local sub-machine can independently solve its allotted subproblem. Then, the central and local sub-machines communicate and coordinate solutions via variable updates. This distributed storage allows for parallel processing and efficient handling of large-scale data. For a more detailed discussion, please refer to Chapters 7 and 8 in [27].
@@
$\min_{\bm{\beta},\bm{r}_{g}}\$$\sum_{g=1}^{G}\sum_{i=1}^{n_{g}}\Big[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\Big]+\lambda\|\bm{w}\odot\bm{\beta}\|_{1},$
s.t. $\bm{r}_{g}=\bm{X}_{g}\bm{\beta},\ g=1,2,\dots,G.$ (4.13)
It is not difficult to verify that (4) and (4.2) are equivalent. The augmented Lagrangian of (4.2) is given by
$L_{\mu}(\bm{\beta},\bm{r}_{g},\bm{u}_{g})$ $=\sum_{g=1}^{G}\sum_{i=1}^{n_{g}}\Big[-S_{i}r_{i}+m\log\left(1+\exp({r_{i}})\right)\Big]$ (4.14)
$+\lambda\|\bm{w}\odot\bm{\beta}\|_{1}-\sum_{g=1}^{G}\bm{u}_{g}^{\top}(\bm{X}_{g}\bm{\beta}-\bm{r}_{g})+\frac{\mu}{2}\sum\limits_{g=1}^{G}\|\bm{X}_{g}\bm{\beta}-\bm{r}_{g}\|_{2}^{2},$
where $\bm{u}_{g}\in\mathbb{R}^{n_{g}}$ is the dual variable and $(\bm{u}_{1}^{\top},\bm{u}_{2}^{\top},\dots,\bm{u}_{G}^{\top})^{\top}=\bm{u}$. Similar to (4.3), the update of the variables in the iterative process (with linearization) can be written as
$\left\{\begin{aligned} \bm{\beta}^{k+1}\leftarrow&\mathop{\arg\min}\limits_{\bm{\beta}}\left\{L_{\mu}(\bm{\beta},\bm{r}_{g}^{k},\bm{u}_{g}^{k})+\frac{1}{2}\|\bm{\beta}-\bm{\beta}^{k}\|^{2}_{{\bm{E}}_{1}}\right\};\\ \bm{r}_{g}^{k+1}\leftarrow&\mathop{\arg\min}\limits_{\bm{r}}\left\{L_{\mu}(\bm{\beta}^{k+1},\bm{r}_{g},\bm{u}_{g}^{k})+(c_{0}-o(1))\|\bm{r}_{g}-\bm{r}_{g}^{k}\|^{2}_{\bm{I}_{n_{g}}}\right\},g=1,2,\dots,G;\\ \bm{u}_{g}^{k+1}\leftarrow&\bm{u}_{g}^{k}-\mu(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{r}_{g}^{k+1}),g=1,2,\dots,G.\end{aligned}\right$. (4.15)
@@
Looking back at the equation in (4.7), in a distributed environment, the update of $\bm{\beta}$ encounters two issues. One issue is that the complete dataset $\bm{X}$ is unavailable (since $\bm{X}$ is stored in a distributed manner and we can only access $\bm{X}_{g}$). The other issue is the computation of the value of $\bm{\eta}$ based on $\bm{X}$. Based on the suggestion in [34], we can reformulate the update of $\bm{\beta}$ in (4.6) as
$\bm{\beta}^{k+1}\leftarrow\underset{\bm{\beta}}{\mathrm{argmin}}\left\{\lambda\left\lVert\bm{w}\odot\bm{\beta}\right\rVert_{1}+\frac{\eta}{2}\left\lVert\bm{\beta}-\bm{\beta}^{k}+\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right\rVert_{2}^{2}\right\},$ (4.16)
@@
$\bm{\beta}^{k+1}\leftarrow\text{sign}\left(\bm{\beta}^{k}-\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right)\odot\left(\left|\bm{\beta}^{k}-\sum_{g=1}^{G}\bm{\xi}_{g}^{k}\right|-\frac{\lambda\bm{w}}{\eta}\right)_{+}.$ (4.17)
Therefore, the first issue with the $\bm{\beta}$ update has been resolved. The second issue pertains to the calculation of $\eta$. The calculation of $\eta$ depends on $\bm{X}$. However, each sub-machine only has $\bm{X}_{g}$. According to the suggestion in [34], $\eta_{g}>\|\mu\bm{X}_{g}^{\top}\bm{X}_{g}\|$ for each sub-machine can be calculated first. Since $\bm{X}^{\top}\bm{X}=\sum\limits_{g=1}^{G}\bm{X}_{g}^{\top}\bm{X}_{g}$, we have $\sum\limits_{g=1}^{G}\eta_{g}>\eta$. Therefore, $\sum\limits_{g=1}^{G}\eta_{g}$ can be used to substitute the $\eta$ calculated from the full data. So far, two issues related to the update of $\bm{\beta}$ in distributed environments have been resolved.
Similar to the derivation of the closed form solution shown in (4.10), we can obtain the update rule for $\bm{r}_{g}$ in a distributed environment,
$\bm{r}_{g}^{k+1}\leftarrow\left[\Phi(\bm{r}_{g}^{k})+\mu\bm{I}_{n_{g}}\right]^{-1}\left[\mu(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{u}_{g}^{k}/\mu)+\Phi(\bm{r}_{g}^{k})\bm{r}_{g}^{k}-\phi^{{}^{\prime}}(\bm{r}_{g}^{k})+\bm{S}_{g}\right],$
@@
   Central machine:    1. Receive $\eta_{g}$ (only once) and $\bm{\xi}_{g}^{k}$ transmitted by all sub-machines,   2. Update $\bm{\beta}^{k+1}$ using (4.17),    3. Send $\bm{\beta}^{k+1}$ to the each sub-machines.
   Local machines: for $g=1,2,\dots,G$ (in parallel)    1. Receive $\bm{\beta}^{k+1}$ transmitted by the central machine,    2. Update $\bm{r}^{k+1}$ using (4.2),    3. Update $\bm{u}_{g}^{k+1}\leftarrow\bm{u}_{g}^{k}-\mu(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{r}_{g}^{k+1}),$    4. Compute $\bm{\xi}_{g}^{k+1}=\frac{\mu\bm{X}_{g}^{\top}(\bm{X}_{g}\bm{\beta}^{k+1}-\bm{r}_{g}^{k+1}-\bm{u}_{g}^{k+1}/\mu)}{\eta}$,    5. Send $\bm{\xi}^{k+1}_{g}$ to the central machine.
@@
$\left\{\hat{\bm{\beta}}^{k},\hat{\bm{r}}^{k},\hat{\bm{d}}^{k}\right\}=\left\{\bar{\bm{\beta}}^{k},\bar{\bm{r}}^{k},\bar{\bm{d}}^{k}\right\},\quad\text{for all }k.$ (4.18)
[…]
5 Numerical Experiments
In this section, we conduct comprehensive evaluations of the proposed method, theoretical framework, and algorithms using both synthetic and real-world datasets. All computational experiments were performed in R on a workstation equipped with an AMD Ryzen 9 7950X 16-Core processor (4.50 GHz) and 32 GB RAM. In the process of parameter tuning, we utilize the modified HBIC criterion presented in [34]. Moreover, the termination conditions of the algorithm adhere to those described in [27].
5.1 Synthetic Data Experiment
We illustrate the methodology using the penalized logistic regression model of [8]. In this example, we simulated 100 datasets consisting of $n$ observations from the model $y\sim\text{Binomial}\{p(\bm{x}^{\top}\bm{\beta}^{*})\}$, where $p(u)=\exp(u)/(1+\exp(u))$ and $\bm{\beta}^{*}=(3,0,0,1.5,0,0,7,0,0)$. The components of $\bm{x}$ are standard normal, where the correlation between $x_{i}$ and $x_{j}$ is $\rho=0.75$.
First, we present the performance of the PLR with manual labels for different values of $m$ and $\alpha_{0}$. Figure 2 shows how the coefficient estimation values change with $m$ and $\alpha_{0}$, where the estimations are averaged over 100 simulations. In the three sub-figures in the upper part of Figure 2, $\alpha_{0}$ is set to 1. From these figures, we can observe that as $m$ increases, the results of coefficient estimation get better, especially for the locus with a value of 7. In the three sub-figures in the lower half of Figure 2, $m$ is set to 20. From these figures, we can observe that as $\alpha_{0}$ increases, the coefficient estimation performance improves, particularly at the positions with values of 1.5 and 7.
Figure 2: Schematic diagram of coefficient estimation values changing with $m$ and $\alpha_{0}$
Table 1: Simulated asymptotic relative efficiency (ARE) for different $m,n$ and $\alpha_{0}$.
$m$ $\alpha_{0}$ ARE $n=100$ $n=200$ $n=500$ $n=1000$ $n=2000$
5 1 1.67 1.55 (0.08) 1.61 (0.08) 1.72 (0.08) 1.70 (0.07) 1.68 (0.07)
10 3.67 3.21 (0.16) 3.26 (0.17) 3.35 (0.19) 3.54 (0.18) 3.69 (0.11)
100 4.81 4.27 (0.23) 4.38 (0.24) 4.45 (0.25) 4.61 (0.24) 4.79 (0.25)
1000 4.98 4.33 (0.25) 4.41 (0.26) 4.53 (0.29) 5.13 (0.31) 4.95 (0.28)
10 1 1.82 1.61 (0.09) 1.66 (0.09) 1.69 (0.10) 1.72 (0.11) 1.79 (0.12)
10 5.50 5.09 (0.21) 5.13 (0.22) 5.24 (0.25) 5.31 (0.24) 5.40 (0.26)
100 9.18 8.55 (0.39) 8.62 (0.43) 8.71 (0.44) 8.89 (0.46) 9.11 (0.49)
1000 9.91 8.96 (0.42) 9.08 (0.45) 9.46 (0.48) 9.63 (0.50) 9.75 (0.52)
50 1 1.96 1.73 (0.10) 1.78 (0.11) 1.82 (0.12) 1.89 (0.13) 1.93 (0.14)
10 9.17 8.30 (0.35) 8.42 (0.38) 8.66 (0.41) 8.89 (0.42) 9.08 (0.43)
100 33.67 28.91 (1.37) 29.21 (1.45) 30.24 (1.58) 35.58 (1.85) 34.07 (1.65)
1000 47.67 40.33 (1.72) 42.74 (1.78) 44.17 (1.93) 45.27 (2.04) 48.01 (2.08)
In Table 1, we compare the simulated asymptotic relative efficiency (ARE) for different values of $m$, $n$, and $\alpha_{0}$. The ARE values presented in the table represent the theoretical values. The estimated ARE values shown in the table are the average results of 100 simulations, with the standard deviation provided in parentheses. In this table, as the values of $m$, $n$ and $\alpha_{0}$ increase, the estimated ARE will perform better and better. This is consistent with the conclusion of consistent estimators in statistics and the theoretical results revealed in this paper.
For large-scale data, we set $n$ to be $1,000,000$ and $2,000,000$ respectively, and conduct numerical simulations in a parallel computing environment where $G$ takes the values of $1,5,10$, and $20$. To verify the conclusion of Theorem 4, the algorithms proposed in this paper all use $\eta=\|\mu\bm{X}^{\top}\bm{X}\|+0.01$. Most of the existing parallel algorithms designed for distributed storage data are inspired by [27]. They utilize consensus constraints to attain algorithm parallelism. For more details, refer to [35] and the references cited therein. Here, we will compare the partition insensitive parallel ADMM algorithm (PIPADMM) proposed in this paper with the consensus structure parallel ADMM algorithm (CPADMM) provided in [35] for solving PLR. Table 2 presents the mean results of 100 independent simulations, with the standard deviation in parentheses. The following is an explanation of the various indicators in the table. FP is a false positive (zero coefficient is estimated to be non-zero), FN is a false negative (non-zero coefficient is estimated to be zero), AE represents the absolute error of estimation ($\|\hat{\bm{\beta}}-\bm{\beta}^{*}\|_{1}$), Ite is the number of iterations, and Time represents the computer running time.
One of the most prominent manifestations is that as the number of parallel machines increases, in addition to a significant reduction in the computation time of PIPADMM, all other indicators remain unchanged. This is also what Theorem 4 in this paper refers to as partition insensitivity, that is, regardless of how the samples are partitioned, the solution of each step of the PIPADMM algorithm remains the same. The reduction in computation time is due to the parallel update of $\bm{r}$. In contrast, for the CPADMM algorithm, as the value of $G$ increases, the quality of the solution (FP, AE) deteriorates, and the number of iterations (Ite) increases. Similarly, due to parallel computing, the computation time significantly decreases. The decline in solution quality and the increase in the number of iterations are due to the inherent flaws of the parallel ADMM algorithm with a consensus structure. As the number of machines increases, more consensus constraints are constructed, leading to more variables to be updated in the ADMM algorithm. This results in lower solution accuracy and slower algorithm convergence, which is also observed in [36] when solving linear regression models. In the comparison between the two parallel ADMM algorithms, PIPADMM outperforms in all metrics. Due to its theoretical property of partition insensitivity, PIPADMM will perform even better than CPADMM as the value of $G$ increases. Overall, the experimental results in Table 2 indicate that our proposed PIPADMM algorithm has more stable, accurate, and faster computational performance in computing PLR compared to CPADMM algorithms in parallel computing environments.
Table 2: Comparison results of parallel ADMM algorithms under large-scale data.
$n=1,000,000$ (PIPADMM) $n=1,000,000$ (CPADMM)
G FP FN AE Ite Time FP FN AE Ite Time
1 0 0 0.024(0.005) 92.3(8.52) 20.4(2.57) 0.17 0 0.075(0.012) 121.2(11.3) 27.4(3.61)
5 0 0 0.024(0.005) 92.3(8.52) 5.40(0.85) 0.21 0 0.106(0.020) 167.5(18.9) 7.41(1.26)
10 0 0 0.024(0.005) 92.3(8.52) 2.92(0.50) 0.26 0 0.162(0.041) 205.1(27.9) 4.76(0.85)
20 0 0 0.024(0.005) 92.3(8.52) 1.42(0.27) 0.31 0 0.199(0.066) 226.3(31.8) 2.98(0.51)
$n=2,000,000$ (PIPADMM) $n=2,000,000$ (CPADMM)
G FP FN AE Ite Time FP FN AE Ite Time
1 0 0 0.019(0.003) 97.2(9.09) 41.3(3.98) 0.12 0 0.162(0.023) 145.8(12.7) 59.7(7.92)
5 0 0 0.019(0.003) 97.2(9.09) 10.8(1.05) 0.14 0 0.205(0.034) 159.7(15.2) 38.2(4.76)
10 0 0 0.019(0.003) 97.2(9.09) 6.29(0.53) 0.17 0 0.396(0.045) 190.4(20.5) 22.4(2.97)
20 0 0 0.019(0.003) 97.2(9.09) 3.92(0.29) 0.23 0 0.471(0.052) 213.5(27.2) 16.1(1.72)
5.2 Real Data Experiments